OpenStudy (anonymous):

Associated with each rubber ball is a bounce coefficient b. When the ball is dropped from a height h, it bounces back to a height of b h. Suppose that the ball is dropped from an initial height h, and then is allowed to bounce forever. Use the expansion of 1/(1 - x) in powers of x to help come up with a clean formula that measures, in terms of b and h, the total up-down distance the ball travels in all of its bouncing.

OpenStudy (rogue):

Oh, we did this last Friday! Hold on, let me get out my notes...

OpenStudy (anonymous):

1 - = 1/1- x )hb 2

OpenStudy (rogue):

\[Height_{total} = H\]\[H = h + 2bh + 2b^2 h + 2b^3 h + ...\]\[H = h (1 + 2b + 2b^2 + 2b^3 + ...)\]\[H = h( 1 + 2b (1 + b + b^2 + b^3 +...))\]\[H = h \left[ 1 + 2b (\frac {1}{1-b}) \right] \]And that nicely becomes\[H = h \left[ \frac {1 + b}{1-b} \right]\]Done!

OpenStudy (anonymous):

Thank You!!!!!!

OpenStudy (rogue):

No problem, its a fun little geometric series challenge =)